# Structured Kolmogorov-Arnold Neural ODEs for Interpretable Learning and Symbolic Discovery of Nonlinear Dynamics

### v2

@@ Abstract
Understanding and modeling nonlinear dynamical systems is a fundamental challenge across science and engineering. Deep learning has shown remarkable potential for capturing complex system behavior, yet achieving models that are both accurate and physically interpretable remains difficult. To address this, we propose Structured Kolmogorov-Arnold Neural ODEs (SKANODEs), a framework that integrates structured state-space modeling with Kolmogorov-Arnold Networks (KANs). Within a Neural ODE architecture, SKANODE employs a fully trainable KAN as a universal function approximator to perform virtual sensing, recovering latent states that correspond to interpretable physical quantities such as displacements and velocities. Leveraging KAN’s symbolic regression capability, SKANODE then extracts compact, interpretable expressions for the system’s governing dynamics. Extensive experiments on simulated and real-world systems demonstrate that SKANODE achieves superior predictive accuracy, discovers physics-consistent dynamics, and reveals complex nonlinear behavior. Notably, it identifies hysteretic behavior in an F-16 aircraft and recovers a concise symbolic equation describing this phenomenon. SKANODE thus enables interpretable, data-driven discovery of physically grounded models for complex nonlinear dynamical systems.
@@ 1 Introduction
Contributions: A new modeling framework, termed Structured Kolmogorov-Arnold Neural ODEs (SKANODEs), is proposed for interpretable and data-driven equation discovery of nonlinear dynamical systems, under partial observability. The main contributions of this work include:
  • 1.
    A physically structured Neural ODE architecture, incorporating both a structured state-space and a structured observation model. This enables interpretable latent state learning from indirect sensor measurements (e.g., accelerations), performing virtual sensing without requiring full observation of physical coordinates.
  • 2.
    A dual-stage application of the Kolmogorov-Arnold Network (KAN); first, as a universal function approximator for flexible dynamics learning, and subsequently as a symbolic learner to extract governing equations in closed form.
  • 3.
    A unified and end-to-end trainable pipeline for symbolic equation discovery, effective in noisy, sensor-limited environments, and capable of recovering interpretable, physics-consistent models directly from raw data.
  • 4.
    Extensive experimental validation on both simulated and real-world systems, including the canonical Duffing and Van der Pol oscillators, where SKANODE successfully recovers governing equations consistent with ground truth, and a real-world F-16 aircraft system, where the model identifies hysteretic behavior and derives an interpretable symbolic equation capturing this nonlinear phenomenon.
Through extensive evaluations on both simulated and real-world systems, it is demonstrated that SKANODE achieves significant predictive accuracy while enabling transparent equation-level understanding of nonlinear dynamics. This opens up new possibilities for applying deep learning in scientific modeling, engineering diagnostics, and physics-based decision making.
@@ 2.3 Symbolic Equation Discovery and Kolmogorov–Arnold Networks
More recent techniques such as Sparse Identification of Nonlinear Dynamics (SINDy) [13] and its extensions [14, 15] apply sparse regression over pre-defined function libraries to recover concise governing equations. Despite achieving strong performance in controlled scenarios, SINDy-like methods depend heavily on (i) access to full state variables and their derivatives, (ii) carefully selected candidate functions, and (iii) accurate numerical differentiation—assumptions that are often violated in real-world, noisy or partially observed systems.
To improve scalability and expressiveness, differentiable symbolic regression frameworks such as AI Feynman [49], PySR [50], and neural symbolic regression (e.g., [19, 51, 52]) employ neural-guided search or evolutionary optimization to discover analytic expressions. These methods offer greater automation but retain critical limitations: they often function as post hoc pipelines requiring clean, fully observed data and are rarely integrated directly into architectures like NODEs [48].
@@ 4 Results
For all comparative experiments, care was taken to ensure a consistent training setup across models. The only architectural difference lies in the use of the KAN in SKANODE, whereas ANODE and SONODE adopt MLPs. The architectures of the KAN modules used in SKANODE are visualized in the corresponding experimental subsections. In our implementation, B-splines are adopted as the spline basis, with 5th-order B-splines specifically employed. The activation functions are modeled via B-splines with a residual connection, which is a typical design feature of KANs. Each KAN is initialized with 6 knots (forming 5 intervals), and the knots are adaptively updated during the training process. For MLP-based models, a 5-layer network with 32 hidden units per layer is adopted throughout. Identical data preprocessing, batch sizes, learning rates schedules, and total training epochs were applied across all models to ensure a fair comparison. The experiments are conducted on a local workstation with an NVIDIA GeForce GTX 5080 GPU. Each experiment is run for 2,000 epochs. The data and code used in this paper are publicly available on GitHub at https://github.com/liouvill/SKANODE.
@@ 4.1 Duffing Oscillator
The performance of SKANODE is evaluated and compared against the equivalent state of the art in the field, namely ANODE and SONODE, on this system, as shown in the top-right subfigure of Figure 2. For benchmarking the symbolic learning capability, a numerical baseline is constructed by directly integrating the measured acceleration signals to obtain velocity and displacement, which are then supplied to SINDy to extract a compact symbolic expression. The predicted trajectories from this numerical benchmark are generated by solving the discovered equations and serve to assess the accuracy of the identified dynamics. It is important to note that this numerical baseline serves solely as a reference for benchmarking the symbolic discovery process; it is not a predictor by itself, as it merely integrates the measured signals without learning the underlying dynamics prior to extracting a symbolic form of the governing equations, and its predictions completely depend on the quality of the extracted equation. In contrast, the proposed SKANODE framework functions both as a dynamics learner—capable of end-to-end prediction of system responses—and as a symbolic extractor that uncovers governing equations from the learned latent quantities.
@@
Noise
level
Model
acceleration
displacement velocity
@@
To quantitatively evaluate model performance, mean squared errors (MSE) are computed for both predicted accelerations and inferred latent states against the ground truth. The mean and standard deviation of MSE values over three independent runs are reported in Table 1. All models achieve similar performance in predicting accelerations, while SKANODE consistently shows slightly superior accuracy. More pronounced differences appear in the recovery of latent states: SKANODE achieves the lowest MSE and variance, indicating its ability to extract physically interpretable dynamics reliably across repeated trials. In contrast, ANODE displays high variance and poor latent state interpretability, confirming its tendency to learn arbitrary latent representations. SONODE produces more stable but still non-interpretable latent trajectories, as its structure does not explicitly align with the true physical coordinates. Furthermore, a structured state-space Neural ODE using a conventional multilayer perceptron (MLP) in place of KAN is also evaluated—referred to as S3NODE in Table 1—implemented as a 5-layer MLP with 32 hidden units per layer. These results indicate that the advantage of SKANODE stems both from its structured inductive biases in the latent and observation models and from the symbolic discovery capability of KAN, which improves interpretability and predictive accuracy by uncovering governing equations and enhancing the estimation of system quantities.
@@
Noise
level
Model
acceleration
displacement velocity
@@
To complement MSE and provide a more diagnostic perspective, the Structural Similarity Index (SSIM) is also reported. SSIM captures the perceptual similarity between predicted and ground-truth trajectories. Results across different levels of observation noise are summarized in Table 2. The SSIM analysis reinforces the conclusions drawn from MSE: SKANODE consistently attains near-perfect scores (>0.99>0.99 across all states and noise levels), indicating that its predictions are not only numerically accurate but also structurally faithful. S3NODE also demonstrates strong performance but remains consistently below SKANODE, while SONODE yields moderate similarity and ANODE fails to converge under noisy conditions.
@@ 4.2 Van der Pol Oscillator
Figure 3: Results on the Van der Pol oscillator. Top left: Symbolic governing equation discovered by $\text{KAN}_{\text{symbolic}}$, compared against the numerical baseline. The identified node receiving displacement input $x$ correctly captures the expected quadratic nonlinearity. Top right: Predicted system observables (accelerations) obtained using the SKANODE framework. SKANODE accurately reconstructs accelerations directly from the inferred latent dynamics, without the need for a separate observation model. Bottom: Inferred latent state variables compared to ground truth. The latent states recovered by SKANODE correspond closely to physically meaningful displacement and velocity trajectories, unlike ANODE and SONODE, whose latent states remain abstract representations without clear physical interpretation.
The performance of SKANODE, ANODE, and SONODE is compared on this system. The symbolic governing equation extracted by $\text{KAN}_{\text{symbolic}}$ within SKANODE is shown in Figure 3, where the identified expression accurately reflects the true system dynamics, outperforming the numerical baseline. The symbolic network clearly reveals a distinct quadratic node that receives the displacement input $x$ and multiplies with the velocity input $\dot{x}$, consistent with the nonlinear damping characteristic of the Van der Pol oscillator. The corresponding predicted accelerations and recovered latent states are also presented. SKANODE accurately reconstructs both the observable accelerations and the latent states, recovering physically meaningful displacement and velocity trajectories. In contrast, ANODE fails to capture the system dynamics entirely, producing large deviations even in reconstructing the measured accelerations. SONODE performs better than ANODE in predicting the observables but still exhibits notable discrepancies and does not recover physically interpretable latent states. Its latent representations remain abstract, as the model lacks explicit physical structure. These results highlight the advantage of SKANODE in simultaneously achieving high predictive accuracy and interpretable latent dynamics through its structured state-space formulation and symbolic discovery capability.
@@ 4.3 F-16 Aircraft
The last example expands to a real-world complex system. The study presented by [62] details the experimental data collected from a full-scale F-16 aircraft during a ground vibration test master class. The dataset is publicly available at https://data.4tu.nl/articles/_/12954911. To simulate real-world conditions, two dummy payloads were mounted on the wing tips of the aircraft. The setup includes 145 acceleration sensors distributed across the aircraft, under excitation of a shaker installed beneath the right wing to apply multisine excitations over a frequency range of 2 to 15 Hz. Data from three specific sensors were made public: one at the excitation point, one on the right wing near the nonlinear interface of interest, and one on the payload adjacent to the same interface. These measurements were taken at a sampling frequency of 400 Hz. The interfaces, consisting of T-shaped connectors on the payload side, slid through a rail attached to the wing side, were identified by preliminary investigation as primary sources of nonlinearity in the aircraft’s structural dynamics, particularly at the right-wing-to-payload interface.
@@
Figure 5: Results on the F-16 aircraft. Top left: Symbolic governing equation discovered by $\text{KAN}_{\text{symbolic}}$. Top right: Predicted system observables (accelerations). Bottom: Inferred latent states visualized in phase portraits. The phase plot obtained by SKANODE exhibits distinct closed-loop patterns characteristic of hysteretic behavior, whereas the phase plot from ANODE yields abstract latent trajectories that lack clear physical interpretation.
The predicted vibration responses are shown in Figure 5. SKANODE achieves comparable accuracy to ANODE in reconstructing the observed accelerations. However, the phase portraits of the inferred latent states reveal that SKANODE produces a much more structured and physically meaningful phase space, exhibiting characteristic loop patterns indicative of hysteretic behavior. This observation is consistent with the findings of [63], confirming the presence of nonlinear hysteresis at the interface. Such interpretability offers important practical value: identifying hysteretic signatures in the dynamics can provide insights into localized energy dissipation mechanisms and evolving structural degradation. This information may support engineers in diagnosing potential fatigue-related issues, monitoring wear at critical joints, and informing predictive maintenance strategies in complex aerospace structures.
@@
$\dot{\mathbf{z}}=\begin{bmatrix}\dot{x}\\ \dot{v}\\ \dot{a}_{h}\end{bmatrix}=\begin{bmatrix}v\\ -(\alpha x+(1-\alpha)a_{h})+\frac{1}{m}u\\ h_{\theta}(v,a_{h})\end{bmatrix}$,
@@
In comparison, the numerical baseline using SINDy fails to identify an equation incorporating the internal hysteretic mechanism. This is because it directly integrates the acceleration signals and seeks an explicit symbolic relation between acceleration, displacement, and velocity, without modeling any latent hysteresis effects. The identified equation from the numerical baseline, as given by
$a=0.0846-9.2480x-11.0607v+1.2781u+720.9876xv-12.2085x^{2}v,$
can reconstruct the observed accelerations to some extent but does not capture the true underlying system dynamics or reveal the implicit hysteretic variable.
Figure 6: Prediction error distribution for the F-16 acceleration response. SKANODE errors are more tightly clustered around zero and exhibit a distribution closer to normality compared with the deep learning baseline (ANODE) and the numerical SINDy baseline, indicating higher accuracy and more stable predictions.
Table 3: Model performance on F-16 acceleration prediction measured in MSE and SSIM. SKANODE outperforms both the deep learning baseline (ANODE) and the numerical baseline.
@@
Numerical 0.0442 0.5868
ANODE 0.0456 0.5684
SKANODE 0.0343 0.6185
We further assess the model performance using both MSE and SSIM metrics. As reported in Table 3, the proposed SKANODE consistently outperforms the deep learning baseline (ANODE) as well as the numerical SINDy baseline in predicting accelerations. The distribution of prediction errors is revealed in Figure 6, where SKANODE exhibits errors that are more tightly clustered around zero and exhibit a distribution closer to normality compared with the baselines. This indicates that SKANODE not only achieves higher accuracy but also provides more stable and reliable predictions. Overall, this example underscores the capability of SKANODE to effectively capture complex, real-world nonlinear dynamics through the structured latent state design and interpretable symbolic equation discovery.

### v3

@@ Abstract
Understanding and modeling nonlinear dynamical systems is a fundamental challenge across science and engineering. Deep learning has shown remarkable potential for capturing complex system behavior, yet achieving models that are both accurate and physically interpretable remains difficult. To address this, we propose Structured Kolmogorov-Arnold Neural ODEs (SKANODEs), a framework that integrates structured state-space modeling with Kolmogorov-Arnold Networks (KANs). Within a Neural ODE architecture, SKANODE employs a fully trainable KAN as a universal function approximator to perform virtual sensing, recovering latent states that correspond to interpretable physical quantities such as displacements and velocities. Leveraging KAN’s symbolic regression capability, SKANODE then extracts compact, interpretable expressions for the system’s governing dynamics. Experiments on two canonical nonlinear oscillators and a real-world F-16 ground vibration dataset demonstrate that SKANODE reliably recovers physically meaningful latent displacement and velocity trajectories from acceleration measurements, identifies the correct governing nonlinearities—including the cubic stiffness in the Duffing oscillator and the nonlinear damping structure in the Van der Pol oscillator—and reveals hysteretic signatures in the F-16 interface dynamics through structured latent phase portraits and an interpretable symbolic model. Across all three cases, SKANODE provides more accurate and robust predictions than black-box NODE baselines and classical ARX and NARX identification, while producing equation-level descriptions of the learned nonlinear dynamics.
@@ 1 Introduction
The resulting framework supports both prediction and equation-level interpretation under partial observability. Through extensive evaluations on simulated benchmark systems and a real-world F-16 ground vibration dataset, it is demonstrated that SKANODE achieves strong predictive accuracy while providing transparent, interpretable descriptions of nonlinear dynamics. This capability broadens the practical use of deep learning for scientific modeling, engineering diagnostics, and physics-based decision making.
@@ 2.3 Symbolic Equation Discovery and Kolmogorov–Arnold Networks
More recent techniques such as Sparse Identification of Nonlinear Dynamics (SINDy) [5] and its extensions [46, 49, 41] apply sparse regression over pre-defined function libraries to recover concise governing equations. Despite achieving strong performance in controlled scenarios, SINDy-like methods depend heavily on (i) access to full state variables and their derivatives, (ii) carefully selected candidate functions, and (iii) accurate numerical differentiation—assumptions that are often violated in real-world, noisy or partially observed systems.
To improve scalability and expressiveness, differentiable symbolic regression frameworks such as AI Feynman [58], PySR [8], and neural symbolic regression (e.g., [23, 25, 22]) employ neural-guided search or evolutionary optimization to discover analytic expressions. These methods offer greater automation but retain critical limitations: they often function as post-hoc pipelines requiring clean, fully observed data and are rarely integrated directly into architectures like NODEs [64].
@@ 4 Results
For all comparative experiments, care was taken to ensure a consistent training setup across models. The only architectural difference lies in the use of the KAN in SKANODE, whereas ANODE and SONODE adopt MLPs. In addition to deep learning baselines, two classical input-output identification models are also included as references: the linear AutoRegressive model with eXogenous inputs (ARX) and its nonlinear extension (NARX). Both methods are fitted directly on the measured acceleration signals, with the same input-output split as SKANODE, and are evaluated on the same test horizon using identical error metrics. For ARX and NARX, only accelerations are predicted directly; the corresponding displacement and velocity trajectories are obtained by numerical time integration of the predicted accelerations using the ground-truth initial conditions.
The architectures of the KAN modules used in SKANODE are visualized in the corresponding experimental subsections. In our implementation, B-splines are adopted as the spline basis, with 5th-order B-splines specifically employed. The activation functions are modeled via B-splines with a residual connection, which is a typical design feature of KANs. Each KAN is initialized with 6 knots (forming 5 intervals), and the knots are adaptively updated during the training process. For MLP-based models, a 5-layer network with 32 hidden units per layer is adopted throughout. Identical data preprocessing, batch sizes, learning rates schedules, and total training epochs were applied across all models to ensure a fair comparison. The experiments are conducted on a local workstation with an NVIDIA GeForce GTX 5080 GPU. Each experiment is run for 1,000 epochs. The data and code used in this paper are publicly available on GitHub at https://github.com/liouvill/SKANODE.
@@ 4.1 Duffing Oscillator
The performance of SKANODE is evaluated and compared against ANODE and SONODE, as well as classical input-output identification baselines ARX and NARX, on this system, as shown in the top-right subfigure of Figure 2. For benchmarking the symbolic learning capability, a numerical baseline is constructed by directly integrating the measured acceleration signals to obtain velocity and displacement, which are then supplied to SINDy to extract a compact symbolic expression. The predicted trajectories from this numerical benchmark are generated by solving the discovered equations and serve to assess the accuracy of the identified dynamics. It is important to note that this numerical baseline serves solely as a reference for benchmarking the symbolic discovery process; it is not a predictor by itself, as it merely integrates the measured signals without learning the underlying dynamics prior to extracting a symbolic form of the governing equations, and its predictions completely depend on the quality of the extracted equation. In contrast, the proposed SKANODE framework functions both as a dynamics learner—capable of end-to-end prediction of system responses—and as a symbolic extractor that uncovers governing equations from the learned latent quantities.
@@
Noise level Model acceleration (measurement) displacement velocity
@@
[…]
0.01 ANODE NaN NaN NaN
[…]
To quantitatively evaluate model performance, mean squared errors (MSE) are computed for both predicted accelerations and inferred latent states against the ground truth. The mean and standard deviation of MSE values over three independent runs are reported in Table 1. All models achieve similar performance in predicting accelerations, while SKANODE consistently shows slightly superior accuracy. More pronounced differences appear in the recovery of latent states: SKANODE achieves the lowest MSE and variance, indicating its ability to extract physically interpretable dynamics reliably across repeated trials. In contrast, ANODE displays high variance and poor latent state interpretability, confirming its tendency to learn arbitrary latent representations. SONODE produces more stable but still non-interpretable latent trajectories, as its structure does not explicitly align with the true physical coordinates. ARX and NARX provide additional reference points from classical system identification. While they can sometimes yield reasonable acceleration predictions, they do not impose physically meaningful state structure and therefore do not support reliable displacement and velocity recovery, especially under measurement noise. Consistent with this limitation, ARX/NARX exhibit substantially larger errors for latent state reconstruction. Furthermore, a structured state-space Neural ODE using a conventional multilayer perceptron (MLP) in place of KAN is also evaluated—referred to as S3NODE in Table 1—implemented as a 5-layer MLP with 32 hidden units per layer. These results indicate that the advantage of SKANODE stems both from its structured inductive biases in the latent and observation models and from the symbolic discovery capability of KAN, which improves interpretability and predictive accuracy by uncovering governing equations and enhancing the estimation of system quantities.
@@ 0.01 ANODE NaN NaN NaN
Noise level Model acceleration (measurement) displacement velocity
@@
To complement MSE and provide a more diagnostic perspective, the Structural Similarity Index (SSIM) is also reported. SSIM captures the perceptual similarity between predicted and ground-truth trajectories. Results across different levels of observation noise are summarized in Table 2. The SSIM analysis reinforces the conclusions drawn from MSE: SKANODE consistently attains highest scores (>0.99>0.99 across all states and noise levels), indicating that its predictions are not only numerically accurate but also structurally faithful. S3NODE also demonstrates strong performance but remains consistently below SKANODE, while SONODE yields moderate similarity and ANODE fails to converge under noisy conditions.
Figure 3: Downstream control study on Duffing regulation, comparing PD (no inversion), feedback linearization using the KAN universal-approximator with Newton inversion (KAN FBL), and feedback linearization using the extracted closed-form KAN symbolic model (KAN symbolic FBL). Top: state regulation trajectories for displacement $x$ (left) and velocity $v$ (right). Bottom left: control effort $u(t)$. Bottom right: box plot of per-step controller computation time (log scale). The symbolic feedback-linearization controller achieves regulation performance comparable to Newton-based inversion while substantially reducing online computation cost.
To further substantiate that the symbolic representation obtained via the proposed scheme, $\text{KAN}_{\text{symbolic}}$, provides operational value beyond post-hoc interpretability, we introduce an additional downstream control study on the Duffing oscillator. The extracted closed-form governing dynamics are first rewritten in control-affine form and then directly embedded into a classical feedback-linearization (FBL) controller.
The regulation task is evaluated under observation noise with standard deviation $\sigma=0.005$, thereby reflecting realistic sensing conditions. Three control strategies are compared: (i) feedback linearization using the extracted closed-form symbolic model (KAN symbolic FBL), (ii) feedback linearization using the KAN universal approximator, which requires online Newton-based inversion at each time step (KAN FBL), and (iii) a baseline proportional–derivative (PD) controller.
[…]
@@ 4.2 Van der Pol Oscillator
Figure 4: Results on the Van der Pol oscillator. NARX trajectories are omitted for visual clarity. Top left: Symbolic governing equation discovered by $\text{KAN}_{\text{symbolic}}$, compared against the numerical baseline. The identified node receiving displacement input $x$ correctly captures the expected quadratic nonlinearity. Top right: Predicted system observables (accelerations) obtained using the SKANODE framework. SKANODE accurately reconstructs accelerations directly from the inferred latent dynamics, without the need for a separate observation model. Bottom: Inferred latent state variables compared to ground truth. The latent states recovered by SKANODE correspond closely to physically meaningful displacement and velocity trajectories, unlike ANODE and SONODE, whose latent states remain abstract representations without clear physical interpretation.
Table 3: Model performance on Van der Pol (VDP) acceleration prediction measured in MSE and SSIM. SKANODE outperforms both deep learning and numerical baselines.
Model MSE SSIM
Numerical 1.0247 0.6783
ARX 3.9180 0.0307
NARX 125.8891 0.0045
ANODE 21.6218 0.1612
SONODE 3.6513 0.1502
SKANODE 0.2317 0.8764
The performance of SKANODE, ANODE, and SONODE, as well as classical identification baselines ARX and NARX, is summarized in Table 3. The symbolic governing equation extracted by $\text{KAN}_{\text{symbolic}}$ within SKANODE is shown in Figure 4, where the identified expression accurately reflects the true system dynamics, outperforming the numerical baseline. The symbolic network clearly reveals a distinct quadratic node that receives the displacement input $x$ and multiplies with the velocity input $\dot{x}$, consistent with the nonlinear damping characteristic of the Van der Pol oscillator. The corresponding predicted accelerations and recovered latent states are also presented. SKANODE accurately reconstructs both the observable accelerations and the latent states, recovering physically meaningful displacement and velocity trajectories. In contrast, ANODE fails to capture the system dynamics entirely, producing large deviations even in reconstructing the measured accelerations. SONODE performs better than ANODE in predicting the observables but still exhibits notable discrepancies and does not recover physically interpretable latent states. Its latent representations remain abstract, as the model lacks explicit physical structure. These results highlight the advantage of SKANODE in simultaneously achieving high predictive accuracy and interpretable latent dynamics through its structured state-space formulation and symbolic discovery capability.
@@ 4.3 F-16 Aircraft
The last example expands to a real-world complex system. The study presented by [37] details the experimental data collected from a full-scale F-16 aircraft during a ground vibration test master class. The dataset is publicly available at https://data.4tu.nl/articles/_/12954911. To simulate real-world conditions, two dummy payloads were mounted on the wing tips of the aircraft. The setup includes 145 acceleration sensors distributed across the aircraft, under excitation of a shaker installed beneath the right wing to apply sine-sweep excitations over a frequency range of 2 to 15 Hz. Data from three specific sensors were made public: one at the excitation point, one on the right wing near the nonlinear interface of interest, and one on the payload adjacent to the same interface. These measurements were taken at a sampling frequency of 400 Hz. In this work, the highest excitation level in this public dataset (95.6 N input amplitude) is considered, under which the system nonlinearity is most pronounced. The interfaces, consisting of T-shaped connectors on the payload side, slid through a rail attached to the wing side, were identified by preliminary investigation as primary sources of nonlinearity in the aircraft’s structural dynamics, particularly at the right-wing-to-payload interface.
@@
Figure 6: Results on the F-16 aircraft. Top left: Symbolic governing equation discovered by $\text{KAN}_{\text{symbolic}}$. Top right: Predicted system observables (accelerations), with the red vertical line marking the training horizon. Bottom: Inferred latent states visualized in phase portraits. The phase plot obtained by SKANODE exhibits distinct closed-loop patterns characteristic of hysteretic behavior, whereas the phase plot from ANODE yields abstract latent trajectories that lack clear physical interpretation.
Figure 7: Time-frequency comparison of the F-16 acceleration response using spectrograms. SKANODE more faithfully reproduces the dominant frequency bands and their temporal evolution.
The predicted vibration responses are shown in Figure 6. Trained only on a short initial segment of the response with substantially smaller amplitude than the subsequent prediction horizon, SKANODE yields a predicted trajectory that matches the measured response significantly better than both the numerical baseline and ANODE. Furthermore, the phase portraits of the inferred latent states reveal that SKANODE produces a much more structured and physically meaningful phase space, exhibiting characteristic loop patterns indicative of hysteretic behavior. This observation is consistent with the findings of [9], confirming the presence of nonlinear hysteresis at the interface. Such interpretability offers important practical value: identifying hysteretic signatures in the dynamics can provide insights into localized energy dissipation mechanisms and evolving structural degradation. This information may support engineers in diagnosing potential fatigue-related issues, monitoring wear at critical joints, and informing predictive maintenance strategies in complex aerospace structures.
In addition, a spectrogram comparison is provided to assess whether the learned models preserve the time-frequency content of the nonlinear response. As shown in Figure 7, SKANODE better preserves the dominant frequency ridges and their time-varying evolution, whereas the baselines exhibit noticeable frequency smearing and mismatched band intensities.
@@
$\dot{\mathbf{z}}=\begin{bmatrix}\dot{x}\\ \dot{v}\\ \dot{a}_{h}\end{bmatrix}=\begin{bmatrix}v\\ -kx-\alpha a_{h}-cv+\frac{1}{m}u\\ h_{\theta}(v,a_{h})\end{bmatrix}$,
@@
In comparison, the numerical baseline using SINDy fails to identify an equation incorporating the internal hysteretic mechanism. This is because it directly integrates the acceleration signals and seeks an explicit symbolic relation between acceleration, displacement, and velocity, without modeling any latent hysteresis effects. The identified equation from the numerical baseline becomes unstable and diverges over the prediction horizon, as the identified expression does not capture the true underlying dynamics of the system.
Table 4: Model performance on F-16 acceleration prediction measured in MSE and SSIM. SKANODE outperforms both deep learning and numerical baselines.
@@
Numerical NaN NaN
ARX 0.5131 0.3597
NARX 0.5085 0.3409
ANODE 0.7139 0.3987
SONODE 0.5324 0.3568
SKANODE 0.0022 0.9576
Figure 8: Prediction error distribution for the F-16 acceleration response. SKANODE errors are more tightly clustered around zero compared with the deep learning and numerical baselines, indicating higher accuracy and more reliable predictions.
We further assess the model performance using both MSE and SSIM metrics. As reported in Table 4, the proposed SKANODE consistently outperforms the deep learning as well as the numerical baselines in predicting accelerations. The distribution of prediction errors is revealed in Figure 8, where SKANODE exhibits errors that are more tightly clustered around zero. This indicates that SKANODE achieves higher accuracy and provides more reliable predictions. Overall, this example underscores the capability of SKANODE to effectively capture complex, real-world nonlinear dynamics through the structured latent state design and interpretable symbolic equation discovery.